# arXiv:2602.22477v2[physics.ins-det] 27 Feb 2026

### v2

@@ 2 Background and Related Work
Cherenkov approaches require no scintillating material or wavelength shifter and provide inherent rejection of sub-threshold radiation, simplifying the detection chain relative to scintillator-based activation detectors. Water-flow 16N monitors count $\beta$-induced Cherenkov light for real-time D–T monitoring in steady-state or long-pulse systems [11]. In clear, non-scintillating glasses, activation products can be read directly via Cherenkov emission, with multi-exponential fits separating contributors by half-life; such glass/Cherenkov methods have even been explored for nuclear-detonation forensics [12, 13].
The Cherenkov photon yield in fused silica is roughly two orders of magnitude lower than typical scintillator light output per MeV of electron energy. Despite this, the observed post-shot count rates ($\sim 10^{3}$$10^{4}\,\text{counts/s}$) are well above the statistical requirements for the multi-exponential fit, and the Cherenkov velocity threshold intrinsically rejects low-energy $\gamma$-ray backgrounds that would contribute to a scintillator signal.
Undoped fused silica (SiO2) offers a simpler, inert, non-toxic alternative that integrates activation target and radiator: short-lived 28Al and 16N $\beta^{-}$ decays produced by $^{28}\mathrm{Si(n,p)}{}^{28}\mathrm{Al}$ and $^{16}\mathrm{O(n,p)}{}^{16}\mathrm{N}$ generate Cherenkov light in the same rod. The result is a compact, D–T-selective diagnostic without Be/As hazards or LaBr-based cost/fragility, and with straightforward optical readout.
@@ 3.1 Fused Silica Optics
The sensor is a high-purity fused-silica (SiO2) rod, $6\text{\,}\mathrm{i}\mathrm{n}$ long by $1\text{\,}\mathrm{i}\mathrm{n}$ in diameter, end-polished. One end of the rod is optically coupled to a Hamamatsu H10580 photomultiplier tube (PMT) assembly that includes a partially active voltage divider base for improved saturation characteristics [14]. The coupling was first made with silicone optical grease and later with EPO-TEK 301 clear epoxy for long-term stability [15]. The assembly is made light-tight with a black 3D-printed PLA cover; no scintillator or wavelength shifter is used (Fig. 1).
@@ 3.2 Activation channels
Because both the 16N ($\sim 10.4$ MeV) and 28Al ($\sim 2.86$ MeV) $\beta^{-}$ endpoints far exceed the 186 keV Cherenkov threshold, the vast majority of decay electrons produce Cherenkov light. Although this emission is concentrated in the highly attenuated UV–blue spectrum, higher-energy electrons generate disproportionately more photons per track, resulting in high detection-weighted efficiency.
@@ 3.3 Electronics
Figure 3: (a) ZEUS post-shot count rate $R(t)$; points: $0.1\text{\,}\mathrm{s}$ bins for shot 241022-06; line: best-fit model with individual components (fit starts at $t=$10\text{\,}\mathrm{s}$\$ to exclude PMT/digitizer saturation). (b) 2D histogram of per-event integral (energy proxy) vs time.
@@ 4.1 Binning and model
Events are accumulated for $600\text{\,}\mathrm{s}$\$ after each shot and binned in $0.1\text{\,}\mathrm{s}$ intervals to form a rate history $R(t)$. We fit
@@
with $\tau_{1}=T_{1/2}^{(^{16}\mathrm{N})}/\ln 2=$10.29\text{\,}\mathrm{s}$\$, $\tau_{2}=T_{1/2}^{(^{28}\mathrm{Al})}/\ln 2=$193.4\text{\,}\mathrm{s}$\$, $\tau_{3}$ free (bounded $\geq$400\text{\,}\mathrm{s}$\$), and $C$ constant, similar to that used for other glass-Cherenkov activation analyses [12, 13]. An example of this bin and fit can be found in Figure 3. The free exponential term (bounded $\tau_{3}\geq$400\text{\,}\mathrm{s}$\$) is included to capture longer-lived contributions such as activation of trace impurities in the fused silica, activation of surrounding structural materials (PTFE reflector, PLA housing), or residual PMT phosphorescence. The constant $C$ captures the PMT dark count rate and slow background radiation. For a typical ZEUS D–T shot, the 16N count rate at $t=$10\text{\,}\mathrm{s}$\$ exceeds the fitted background $(C$ plus the slow exponential)) by roughly two orders of magnitude, and the 28Al component remains above background for approximately $400\text{\,}\mathrm{s}$, providing signal-to-background ratios well suited to the fit.
@@ 4.3 Uncertainties
The primary operational role of the fused-silica detector is to provide high-precision relative pulse-to-pulse yield measurements. Consequently, the intrinsic uncertainty of the diagnostic is dominated by counting statistics and the multi-exponential fit procedure, rather than absolute systematic scale.
@@
For absolute yield reporting, the detector relies on cross-calibration constants ($K_{\mathrm{Al}}$, $K_{\mathrm{N}}$) derived from the facility reference. This introduces a systematic uncertainty inherited from the reference diagnostic (e.g., praseodymium calibration, reference geometry, and plasma anisotropy). Because a full systematic error budget for the facility yield was not available, the calibration constants are scaled to the nominal reported yields; thus the uncertainties reported here reflect only the intrinsic statistical and fitting precision of the SiO2 detector, while the absolute systematic uncertainty acts as a global scaling factor on the final fluence.
@@ 5.2 MJOLNIR D–D insensitivity
The detector operated near the MJOLNIR D–D DPF under the same DAQ settings as specified above. By placing the detector closer to the pinch, it was exposed to a nominal D–D fluence comparable to the lower-end ZEUS D–T shots. No count rate increase was observed besides the prompt detection of the shot (Fig. 5). This null result is attributed to neutron energy rather than fluence: the ENDF/B-VII.1 cross sections for $^{28}\mathrm{Si(n,p)}$ and $^{16}\mathrm{O(n,p)}$ are effectively zero at $2.45\text{\,}\mathrm{M}\mathrm{e}\mathrm{V}$ (Fig. 2), and applying the ZEUS-derived calibration constants shows that $14.1\text{\,}\mathrm{M}\mathrm{e}\mathrm{V}$ neutrons at this nominal fluence would produce a detectable signal. Dense plasma focus devices operating with deuterium can produce a secondary 14.1 MeV component via burn-up of $\sim$1 MeV tritons created in the D(d,p)T branch. Triton burn-up is a standard feature of deuterium fusion plasmas, with the DT/DD ratio depending on triton confinement and plasma density [23, 24, 25]. In magnetically confined deuterium plasmas, measured burn-up fractions $Y_{DT}/Y_{DD}$ are typically $10^{-3}$ to $10^{-2}$ [25, 26]. For short-pulse DPF plasmas, significantly smaller fractions are expected; values of order $10^{-3}$ to $10^{-4}$ provide a conservative bound. At MJOLNIR fluence levels ($10^{8}$ to $10^{9}$ n/cm2 at 2.45 MeV), this implies a secondary 14.1 MeV fluence of $\leq 10^{4}$ n/cm2. The null result is therefore consistent with both the detection threshold and the expected secondary D–T yield.
@@ 6.3 Comparison to alternatives
Classical activation foils (In, Cu, Zr, etc.) achieve more robust measurements but require extraction and extended measurements on dedicated setups; the fused silica activation detector is remotely read out and intrinsically D–T selective. Compared to NIF’s RT-NADs (LaBr3[9], the fused silica approach is lower cost and easily deployable. Relative to prompt scintillation detectors, the activation–Cherenkov method integrates fluence into a decay signal that is read out seconds to minutes after the pulse, inherently avoiding the prompt EMI transient associated with pulsed fusion generators (the fit window begins at $t=$10\text{\,}\mathrm{s}$\$; see Section 4.1). Water-based 16N Cherenkov monitors are ideal for steady/long pulses [11]; a solid fused-silica detector is optimized for discrete pulses and requires no fluid handling. Our undoped fused-silica approach leverages multi-exponential Cherenkov counting demonstrated in glass systems [12, 13] while avoiding toxic or scarce activation materials.
@@ 7 Conclusions
A fused-silica activation Cherenkov detector provides a simple, fast, precise, and D–T-selective neutron-yield diagnostic. It delivers pulse-to-pulse yields within $\sim$$3\text{\,}\mathrm{m}\mathrm{i}\mathrm{n}$ using list-mode acquisition and multi-exponential fitting. Measured yields agree with a calibrated silver activation detector, and the system is insensitive to D–D neutrons. Its compactness, simplicity, and low cost make it suitable for scaling to multiple channels or deployment at smaller facilities. Ongoing Polaris operation will expose the detector to higher fluences and allow evaluation of durability and calibration stability against standard Zr foils.

### v3

# 1]organization=Helion Energy Inc., city=Everett, WA,country=USA

@@ 2 Background and Related Work
Cherenkov approaches require no scintillating material or wavelength shifter and provide inherent rejection of sub-threshold radiation, simplifying the detection chain relative to scintillator-based activation detectors. The Cherenkov photon yield in fused silica is roughly three orders of magnitude lower than typical scintillator light output per MeV of electron energy. With the additional advantage that Cherenkov photons being prompt, having no decay time or afterglow, allowing high count rate applications. Water-flow 16N monitors count $\beta$-induced Cherenkov light for real-time D–T monitoring in steady-state or long-pulse systems [11]. In clear, non-scintillating glasses, activation products can be read directly via Cherenkov emission, with multi-exponential fits separating contributors by half-life; such glass/Cherenkov methods have even been explored for nuclear-detonation forensics [12, 13].
Undoped fused silica (SiO2) offers a simpler, inert, non-toxic alternative that integrates activation target and radiator: short-lived 28Al and 16N $\beta^{-}$ decays produced by $^{28}\mathrm{Si(n,p)}{}^{28}\mathrm{Al}$ and $^{16}\mathrm{O(n,p)}{}^{16}\mathrm{N}$ generate Cherenkov light in the same material. The result is a compact, D–T-selective diagnostic without Be/As hazards or LaBr-based cost/fragility, and with straightforward optical readout.
@@ 3.1 Fused Silica Optics
The sensor is a high-purity fused-silica (SiO2) rod, $6\text{\,}\mathrm{i}\mathrm{n}$ long by $1\text{\,}\mathrm{i}\mathrm{n}$ in diameter that has been end-polished and wrapped in eight layers of PTFE reflector. One end of the rod is optically coupled to a Hamamatsu H10580 photomultiplier tube (PMT) assembly that includes a partially active voltage divider base for improved saturation characteristics [14]. The coupling was first made with silicone optical grease and later with EPO-TEK 301 clear epoxy for long-term stability [15]. The assembly is made light-tight with a black 3D-printed PLA cover; no scintillator or wavelength shifter is used (Fig. 1).
@@ 3.2 Activation channels
Because both the 16N ($\sim 10.4$ MeV) and 28Al ($\sim 2.86$ MeV) $\beta^{-}$ endpoints far exceed the 186 keV Cherenkov threshold, the vast majority of decay electrons produce Cherenkov light. Although this emission is concentrated in the UV–blue spectrum, higher-energy electrons generate disproportionately more photons per track, resulting in high detection-weighted efficiency.
@@ 3.3 Electronics
Figure 3: (a) ZEUS post-shot count rate $R(t)$; points: $0.1\text{\,}\mathrm{s}$ bins for shot 241022-06; line: best-fit model with individual components (fit starts at $t=$1\text{\,}\mathrm{s}$\$ to exclude PMT/digitizer saturation). (b) 2D histogram of per-event integral (energy proxy) vs time.
@@ 4.1 Binning and model
Events are accumulated for $600\text{\,}\mathrm{s}$\$ after each shot and binned in $0.1\text{\,}\mathrm{s}$ intervals to form a rate history $R(t)$. For data $1\text{\,}\mathrm{s}$ past the trigger, we fit
@@
with $\tau_{1}=T_{1/2}^{(^{16}\mathrm{N})}/\ln 2=$10.29\text{\,}\mathrm{s}$\$, $\tau_{2}=T_{1/2}^{(^{28}\mathrm{Al})}/\ln 2=$193.4\text{\,}\mathrm{s}$\$, $\tau_{3}$ free (bounded $\geq$400\text{\,}\mathrm{s}$\$), and $C$ constant, similar to that used for other glass-Cherenkov activation analyses [12, 13]. An example of this bin and fit can be found in Figure 3. The free exponential term (bounded $\tau_{3}\geq$400\text{\,}\mathrm{s}$\$) is included to capture longer-lived contributions such as activation of trace impurities in the fused silica, activation of surrounding structural materials (PTFE reflector, PLA housing), or residual PMT phosphorescence. The constant $C$ captures the PMT dark count rate and slow background radiation. For a typical ZEUS D–T shot, the 16N count rate at $t=$1\text{\,}\mathrm{s}$\$ exceeds the fitted background $(C$ plus the slow exponential)) by roughly two orders of magnitude, and the 28Al component remains above background for approximately $400\text{\,}\mathrm{s}$, providing signal-to-background ratios well suited to the fit.
@@ 4.3 Uncertainties
The fused-silica detector is operated as a precision relative-yield monitor. Accordingly, the intrinsic uncertainty is dominated by counting statistics and the multi-exponential fit, while the absolute-scale uncertainty is calibration-specific and enters as an overall normalization.
@@
For absolute yield reporting, the detector relies on cross-calibration constants ($K_{\mathrm{Al}}$, $K_{\mathrm{N}}$) derived from the ZEUS facility reference or Polaris neutron activation foil measurements. This introduces a systematic uncertainty inherited from the reference diagnostic (e.g., praseodymium calibration, reference geometry, and plasma anisotropy). Because a full systematic error budget for the ZEUS facility yield was not available, the calibration constants are scaled to the nominal reported yields; thus the uncertainties reported here reflect only the intrinsic statistical and fitting precision of the SiO2 detector, while the absolute systematic uncertainty acts as a global scaling factor on the final fluence.
@@ 5.2 MJOLNIR D–D insensitivity
The detector operated near the MJOLNIR D–D DPF under the same DAQ settings as specified above. By placing the detector closer to the pinch, it was exposed to a nominal D–D fluence comparable to the lower-end ZEUS D–T shots. No count rate increase was observed besides the prompt detection of the shot (Fig. 5). This null result is attributed to neutron energy rather than fluence: the ENDF/B-VII.1 cross sections for $^{28}\mathrm{Si(n,p)}$ and $^{16}\mathrm{O(n,p)}$ are effectively zero at $2.45\text{\,}\mathrm{M}\mathrm{e}\mathrm{V}$ (Fig. 2), and applying the ZEUS-derived calibration constants shows that $14.1\text{\,}\mathrm{M}\mathrm{e}\mathrm{V}$ neutrons at this nominal fluence would produce a detectable signal. Dense plasma focus devices operating with deuterium produce secondary 14.1 MeV neutrons via burn-up of approximately 1 MeV tritons from the D(d,p)T reaction branch. Triton burn-up is a standard feature of deuterium fusion plasmas and produces a secondary neutron population whose magnitude depends on triton confinement and plasma density [23, 24, 25]. In magnetically confined plasmas, $Y_{DT}/Y_{DD}$ can reach $10^{-3}$$10^{-2}$. In DPF devices, however, the plasma is short-lived, spatially small, and weakly confining for 1 MeV tritons, so escape losses are expected to strongly suppress burn-up. The resulting secondary 14.1 MeV yield should therefore be well below $10^{-3}$ of the primary D–D yield and does not impose a practical lower bound on the diagnostic sensitivity considered here.
@@ 6.3 Comparison to alternatives
Classical activation foils (In, Cu, Zr, etc.) achieve more robust measurements but require extraction and extended measurements on dedicated setups; the fused silica activation detector is remotely read out and intrinsically D–T selective. Compared to NIF’s RT-NADs (LaBr3[9], the fused silica approach is lower cost and easily deployable. Relative to prompt scintillation detectors, the activation–Cherenkov method integrates fluence into a decay signal that is read out seconds to minutes after the pulse, inherently avoiding the prompt EMI transient associated with pulsed fusion generators (the fit window begins at $t=$1\text{\,}\mathrm{s}$\$; see Section 4.1). Water-based 16N Cherenkov monitors are ideal for steady/long pulses [11]; a solid fused-silica detector is optimized for discrete pulses and requires no fluid handling. Our undoped fused-silica approach leverages multi-exponential Cherenkov counting demonstrated in glass systems [12, 13] while avoiding toxic or scarce activation materials.
@@ 7 Conclusions
A fused-silica activation Cherenkov detector provides a simple, fast, precise, and D–T-selective relative neutron-yield diagnostic. It delivers pulse-to-pulse relative yields within $\sim$$3\text{\,}\mathrm{m}\mathrm{i}\mathrm{n}$ using list-mode acquisition and multi-exponential fitting. Measured relative yields agree with a calibrated silver activation detector, and the system is insensitive to D–D neutrons. Its compactness, simplicity, and low cost make it suitable for scaling to multiple channels or deployment at smaller facilities. Ongoing Polaris operation will expose the detector to higher fluences and allow evaluation of durability and calibration stability against standard Zr foils.